OpenStudy (anonymous):

Simplify the following expression: 8 squared divided by the square root of 16 plus five squared minus eight.

jimthompson5910 (jim_thompson5910):

I'm assuming it's \[\large \frac{8^2}{\sqrt{16}}+5-8\] \[\large \frac{8^2}{\sqrt{16}}+5-8\] \[\large \frac{64}{4}+5-8\] \[\large 16+5-8\] \[\large 13\] So \[\large \frac{8^2}{\sqrt{16}}+5-8=13\]

OpenStudy (anonymous):

yes but the 5 is squared

jimthompson5910 (jim_thompson5910):

oh right, missed that, one sec

jimthompson5910 (jim_thompson5910):

\[\large \frac{8^2}{\sqrt{16}}+5^2-8\] \[\large \frac{64}{4}+25-8\] \[\large 16+25-8\] \[\large 33\] So \[\large \frac{8^2}{\sqrt{16}}+5^2-8=33\]

OpenStudy (anonymous):

THANKS A TONN :D

jimthompson5910 (jim_thompson5910):

np